How would you solve problems like:
x^2-25=0 x=....
x^2-49=0 x=....
Thank you

Answer by bsmath(21)

How would you solve problems like:
x^2-25=0 x=....
x^2-49=0 x=....

You have two quadratic equations. Quadratic equations are equations of degree
2. The degree is the highest exponent of the unknown. Your equations are a
special form because they are x^2 minus a square. So all you have to do is
find the square root, then the product of x minus the square root and x plus
the square root will be the answer. We have, for x^2-25, since the square root
of 25 is 5, (x-5)(x+5)=x^2-25. For x^2-49, since the square root of 49 is 7,
we have (x-7)(x+7)=x^2-49.